ROTATIONALLY RESOLVED IR SPECTRA OF LiD_{2}^{+} AND LiH_{2}^{+}

\underline{CORINNA EMMELUTH}, CHRISTOPHER D. THOMPSON, BERWYCK L. J. POAD, and EVAN J. BIESKE, Laser Spectroscopy Group, School of Chemistry, University of Melbourne, Parkville, VIC 3010 Australia; GARY H. WEDDLE, Fairfield University, Department of Chemistry, Fairfield, CT 06430 USA.

Photodissociation infrared spectra of mass selected LiD_{2}^{+} and LiH_{2}^{+} have been obtained in the D-D and H-H stretch region respectively, using a tandem mass spectrometer and detecting the Li^{+} loss channel. For the first time rotationally resolved spectra of these complexes are available providing structural parameters and allowing direct comparison with theoretical data ^{a}. For LiD_{2}^{+} around 100 lines of the K_{a}=0←K_{a}=0, K_{a}=1←K_{a}=1, and K_{a}=2←K_{a}=2 parallel transitions were fitted to a Watson A-reduced Hamiltonian. The analysis of the spectrum was supported by quantum chemical calculations using the program TRIATOM ^{b} and the potential from Gianturco et al. ^{c}. The LiD_{2}^{+} complex was found to have T-shaped structure in agreement with theoretical predictions. An analogous analysis has been performed for the LiH_{2}^{+} spectrum which differs from the LiD_{2}^{+} spectrum due to larger rotational constants and an altered ortho:para ratio.

aV. P. Bulychev, K. M. Bulanin, and M. O. Bulanin, Mol. Spec. 96, 2004, 205-216

bJ. Tennyson, S. Miller, and C. R. Le Sueur, Comp. Phys. Comm. 75, 1993, 339-364.

cR. Martinazzo, G. F. Tantardini, E. Bodo, and F. A. Gianturco, J. Chem. Phys. 119, 2003, 11241-11248